# TaSL: Task Skill Localization and Consolidation for Language Model Continual Learning

### v1

@@ VI-A1 Datasets
Additionally, we employ the Long Sequence Benchmark [18], a continual learning benchmark consisting of 15 classification datasets. In line with [19], we randomly select 1,000 samples for training each task and reserve 500 samples per class for validation and testing. We explore two different task orders for each benchmark. Please refer to the Appendix for more details about the tasks and orders.
TABLE I: The overall results on two continual learning benchmarks with T5-large model. All results are averaged over two different orders of each benchmark. Our rehearsal-free TaSL outperforms the previous best method, O-LoRA, by achieving an average 2.3% absolute improvement on AP, an average 2.8% absolute improvement on FWT, and an 4.8% absolute increase in BWT.
@@ VI-A3 Baselines
We evaluate TaSL against the following PEFT-based continual learning baseline methods: SeqLoRA: sequentially trains the LoRA on the task orders. IncLoRA: incremental learning of new LoRA parameters on a sequential series of tasks. Replay: replays real samples from old tasks when learning new tasks to avoid forgetting. EWC [46]: finetune the model with a regularization loss that prevents updating parameters that could interfere with previously learned tasks. L2P [49]: uses the input to dynamically select and update prompts from a fixed prompt pool. LFPT5 [50]: continuously trains a soft prompt for each task with generative replay and an auxiliary loss. Prog-Prompt [18]: sequentially concatenates previous learned prompts to the current one during the training and testing time. O-LoRA [19]: learns tasks in different LoRA subspaces that are kept orthogonal to each other and sums all LoRA weights up at testing time. SAPT [16]: leverages pseudo samples and a shared attention framework to align PEFT block learning and selection.
Table II compares our TaSL with these baselines, highlighting three distinct advantages: data privacy-friendliness, model parameter-friendliness, and generalization-friendliness. To clarify the specific localization and consolidation techniques used in TaSL and its variants: TaSL: our base framework utilizes original components, including matrix-level skill units, first-order gradient-based metric for skill localization (Eq. 8), and static weighted consolidation (Eq. 15). TasLoRA: as the latest extension of TaSL, it employs LoRA-tailored skill units, the second-order gradient approximation metric for skill localization (Eq. 9), and adaptive weighted consolidation (Eq. 17). TasL-M and TasLoRA-M: these variants combine TaSL or TasLoRA with memory replay techniques to further enhance performance.
@@ VI-A4 Implementation Details
We utilize two language models adopted by the previous lines of works in CL for NLP: encoder-decoder T5 model [57] and decoder-only LLaMA model [58]. In TaSL, the hyperparameters $\alpha_{1}$ and $\alpha_{2}$ in Eq. (10) and Eq. (11) are set to 0.85. We set $\beta$ in Eq. (13) to 0.7 and $\tau$ in Eq. (16) to 0.15. Following [59], the volume of replay samples is set to 2% of the original training set for all replay-based baseline methods. For different backbones, we utilized the following hyperparameters:
  • T5-base (220M), T5-large (780M), and T5-XL (3B): Training was conducted with a learning rate of 3e-4, batch size of 8, maximum input length of 512, maximum target length of 128, rank = 8, targeting modules [q, v] and 10 epochs.
  • LLaMA (7B): With a learning rate of 3e-4, batch size of 128, a cutoff length of 512, and 10 epochs. Lora settings were rank = 8, alpha = 16, dropout = 0.05, targeting modules [q_proj, v_proj]. For testing, settings included temperature = 0.02, top_p = 0, top_k = 1, max new tokens = 128.
Experiments are carried out using 4 NVIDIA A100.
@@ VI-B Main Results
Overall CL results of different methods at the same T5-large backbone are summarized in Table I.
Firstly, TaSL and its variants demonstrate superior CL performance through effective knowledge transfer. Compared to previous rehearsal-free methods such as IncLoRA and O-LoRA, TaSL excels in addressing the two critical challenges of CF and KT, achieving the highest AP, FWT, and BWT when learning tasks sequentially.
Moreover, our enhanced version, TasLoRA, shows significant improvements over TaSL. Specifically, TasLoRA achieves average gains of 1.3% on AP and 1.4% on BWT, indicating that our extensions to each component are effective. Notably, even without relying on memory replay, our method nearly matches the performance of SAPT with memory replay on Long Sequence Benchmark, particularly in BWT, with only a slight difference of 0.8% (-2.04% vs. -1.25%).
For memory-based methods, both TaSL-M and TasLoRA outperform the best SAPT methods, further demonstrating the versatility and robustness of our framework. For the convenience of subsequent experimental analysis and comparison, we focus on the performance of TasLoRA, the best-performing memory-free version, for further evaluations.
Secondly, TasLoRA consistently demonstrates superior performance across various backbones. To further validate the effectiveness of our framework, we conducted experiments using a range of parameter-level backbones, as shown in Figure 4, which highlight performance improvements with increasing model size. Across different backbones, our method consistently outperforms traditional approaches. For example, in T5-xl, TasLoRA significantly increases the AP metric from 28.6% to 33.3%, and achieves substantial improvements in both FWT and BWT metrics, rising from 11.1% to 16.1% and improving from -20.6% to -14.0%, respectively. These results further validate the robust generalization ability of our proposed framework.
@@
Figure 4: Performance of TaSL w/ different backbones.
Thirdly, adaptive skill consolidation effectively mitigates catastrophic forgetting. To rigorously assess our model’s effectiveness in countering forgetting, we analyzed its performance on the initial task after training on subsequent tasks using the Long Sequence Benchmark. Figure 5 shows that our TasLoRA method significantly slows the forgetting rate, with an average performance decrease of only 11% after training on the final task. In stark contrast, vanilla backbones exhibit a substantial average performance drop of 20%, underscoring our method’s superior ability to mitigate forgetting.
Figure 5: Performance trajectory of Task 1 on Long Sequence Benchmark during the CL process.
@@ VI-C Performance on Unseen Tasks
Following previous work [16], we further select three tasks from each one of the above task categories to assess the TaSL’s cross-task generalization ability. This is also a crucial dimension for evaluating CL algorithms. Table III shows the results. T5-ZS represents the zero-shot approaches for task adaptation, respectively. TasLoRA achieved the best performance, which can be attributed to its ability to identify task-specific and shared parameters effectively.
TABLE III: Results on unseen tasks based on the T5-Large backbone model. We report the average Rouge-L of the 3 tasks under each category.
@@ VI-D Visualization of Skill Units
We visualized the distribution of importance scores for the skill units across tasks on T5-large, as shown in Figure 6 and Figure 7, leading to several critical insights:
  • There is a noticeable variation in the importance of skill units for the same task, with important skill units making up only a small fraction of all trainable LoRA parameters.
  • The distribution of important skill units is task-dependent, indicating both task-shared and specific parameters, confirming TaSL’s validity.
  • For classification tasks, such as those in the Long Sequence Benchmark (Figure 6), the skill units in the encoder, particularly the lower layers closer to the model input, play a more significant role. In contrast, for generative tasks, such as dialogue generation and summarization in the SuperNI benchmark (Figure 7), both the encoder and decoder skill units are important, with both the lower and upper layers of the network being crucial.
  • Within each layer, the importance of the Query (Q) matrices in the attention mechanism consistently surpasses that of the Value (V) matrices.
Figure 6: Visualization of importance scores for skill units across different tasks on T5-large for the Long Sequence Benchmark.
Figure 7: Visualization of importance scores for skill units across different tasks on T5-large for the SuperNI Benchmark.
VI-E Ablation Study
In this section, we evaluate the impact of importance-aware skill localization, fine-grained skill consolidation, and hyperparameter sensitivity, with the results discussed below.
VI-E1 Effect of the proposed importance metric in skill localization
Our method calculates importance scores using Eq. (6). As shown in Table IV, we explore alternative importance scoring approaches: (i) to test the effectiveness of using moving averages on trajectory gradients, we modify $s(\cdot)$ in Eq. (6) to include only sensitivity, as described in Eq. (9); and (ii) to assess the validity of our proposed approximate second-order gradient importance metric in Eq. (6), we use a first-order Taylor expansion to approximate the importance $I$ by using Eq. (8). The results indicate that using moving averages for importance scoring outperforms the alternatives, with the other two variants leading to performance decreases of up to 3.3%, 2.5%, and 4.2% across the three metrics. This underscores the value of accurate skill localization in enhancing model performance.
TABLE IV: Ablation study. Evaluating the impact of importance scoring variations on skill localization.
@@ VI-E2 Effect of Adaptive Skill Consolidation
We compared our fine-grained skill unit-level adaptive LoRA averaging mechanism against two coarse-grained strategies: (i) Weight-Ensemble, which uses a global weight to uniformly average the LoRA parameters, i.e., Eq. (14), and (ii) Exponential Moving Average (EMA) [60], which applies a running average of parameters at each fine-tuning iteration.
For Table V, Weight-Ensemble significantly improves upon the vanilla model, highlighting the benefits of coarse-grained averaging for continual learning. EMA generally outperforms Weight-Ensemble but falls short of our fine-grained approach due to its overuse of averaging, with frequent parameter adjustments within the same task potentially leading to suboptimal outcomes. Our method, which averages weights only after each task, enhances computational efficiency.
TABLE V: Ablation study. Comparing coarse- and fine-grained model averaging methods on skill consolidation.
@@
VI-E3 Sensitivity Analysis for Hyperparameters
The proposed framework incorporates three key hyperparameters, including the $\alpha$ for computing importance scores in Eq. (10) and Eq. (11), the $\beta$ for calculating cumulative importance scores in Eq. (13), and the $\tau$ for performing weighted averaging within skill units as outlined in Eq. (16). Our analysis aims to assess the impact of varying these hyperparameters on our method’s performance, testing on the T5-large backbone model.
As evidenced in Table VI, we determine that the optimal setting for $\alpha$ is 0.55. An $\alpha$ value too low results in a performance decline, indicating that the calculated importance scores are not sufficiently accurate. Furthermore, as depicted in Table VII, we also find that $\beta$ values within a normal range do not significantly affect performance. However, excessively high or low values for $\beta$ may skew the model towards favoring either past or current task knowledge, thereby disrupting the desired balance. Nonetheless, the model’s performance remains relatively stable across most conditions, indicating a low sensitivity to hyperparameter variations.
The VIII below shows the model’s performance under different temperature coefficients, $\tau$. It can be seen that setting a higher $\tau$ (higher than 0.5) smooths the weights, thereby increasing the weight of unimportant skill units, which reduces the model’s effectiveness and may contaminate historical knowledge, leading to forgetting. On the other hand, setting a lower $\tau$ (less than 0.1) may cause larger differences in weight coefficients, thereby limiting KT between different tasks.
TABLE VI: Performance comparisons of TaSL equipped with different $\alpha$ on Long Sequence Benchmark at task order 1.
$\alpha_{1}$,$\alpha_{2}$ AP FWT BWT
Vanilla T5-large 51.6 5.6 -29.7
0.15 71.8 13.7 -3.7
0.35 71.2 14.1 -4.3
0.55 72.8 12.6 -2.5
0.85 70.7 12.9 -3.3
0.95 71.7 14.0 -3.6
TABLE VII: Performance comparisons of TaSL equipped with different $\beta$ on Long Sequence Benchmark at task order 1.
$\beta$ AP FWT BWT
Vanilla T5-large 51.6 5.6 -29.7
0.1 73.8 13.6 -3.7
0.3 73.5 12.4 -3.4
0.5 74.3 13.5 -2.2
0.7 72.7 12.9 -2.3
0.9 60.2 14.2 -5.0
TABLE VIII: Performance comparisons of TaSL (using T5-large as the backbone) equipped with different $\tau$ on Long Sequence Benchmark at task order 1.
temperature coefficient $\tau$ AP FWT BWT
0.05 69.0 14.3 -3.4
0.15 73.4 12.8 -2.1
0.25 72.2 13.2 -2.5
0.35 72.1 13.6 -2.1
0.45 72.7 12.5 -3.0
0.55 70.9 11.6 -4.2
@@ VII Conclusion
In this paper, we introduce a novel Task Skill Localization and Consolidation (TaSL) framework for language model continual learning, facilitating effective knowledge transfer across tasks. Our framework leverages an importance-aware skill localization and a fine-grained skill consolidation technique to differentiate between task-specific and shared knowledge within each skill unit, thereby mitigating forgetting. TaSL also demonstrates strong generalizability and extensibility, with optimizations to various components and the ability to integrate with memory replay methods for further performance enhancement. Comprehensive experiments showcase our framework’s exceptional ability to balance preserving past knowledge and excelling in new tasks, surpassing previous state-of-the-art methods.
@@
While our method significantly improves efficiency and performance in large language model continual learning, some limitations remain. For instance, the choice of importance thresholds after determining the importance distribution can affect the effectiveness of subsequent consolidation. Dynamically selecting thresholds based on data distribution could lead to more accurate classification of task-shared and task-specific parameters, thereby enhancing performance. Additionally, we could explore merging the skill localization and consolidation stages, allowing for parameter consolidation based on importance during model training. This would enable more flexible adaptation and address scenarios in online continual learning.
[Dataset Details] Table IX & X show details of the datasets we used for our experiments, along with their evaluation metrics. Overall, in SuperNI, we choose 3 tasks from dialogue generation (Dialog), information extraction (IE), question answering (QA), summarization (Sum) and sentiment analysis (SA), respectively.
For the Long Sequence benchmark, this includes five tasks from the standard CL benchmark (AG News, Amazon reviews, Yelp reviews, DBpedia and Yahoo Answers), four from GLUE benchmark (MNLI, QQP, RTE, SST2), five from SuperGLUE benchmark (WiC, CB, COPA, MultiRC, BoolQ), and the IMDB movie reviews dataset. We report 4 different task orders used for our experiments in Table XI.
TABLE IX: The details of 15 datasets in the SuperNI Benchmark [48].
Dataset name Task Metric
1. task639_multi_woz_user_utterance_generation dialogue generation Rouge-L
2. task1590_diplomacy_text_generation dialogue generation Rouge-L
3. task1729_personachat_generate_next dialogue generation Rouge-L
4. task181_outcome_extraction information extraction Rouge-L
5. task748_glucose_reverse_cause_event_detection information extraction Rouge-L
6. task1510_evalution_relation_extraction information extraction Rouge-L
7. task002_quoref_answer_generation question answering Rouge-L
8. task073_commonsenseqa_answer_generation question answering Rouge-L
9. task591_sciq_answer_generation question answering Rouge-L
10. task511_reddit_tifu_long_text_summarization summarization Rouge-L
11. task1290_xsum_summarization summarization Rouge-L
12. task1572_samsum_summary summarization Rouge-L
13. task363_sst2_polarity_classification sentiment analysis accuracy
14. task875_emotion_classification sentiment analysis accuracy
15. task1687_sentiment140_classification sentiment analysis accuracy
TABLE X: The details of 15 classification datasets in the Long Sequence Benchmark [18]. First five tasks correspond to the standard CL benchmark [61].
Dataset name Category Task Domain Metric
1. Yelp CL Benchmark sentiment analysis Yelp reviews accuracy
2. Amazon CL Benchmark sentiment analysis Amazon reviews accuracy
3. DBpedia CL Benchmark topic classification Wikipedia accuracy
4. Yahoo CL Benchmark topic classification Yahoo Q&A accuracy
5. AG News CL Benchmark topic classification news accuracy
6. MNLI GLUE natural language inference various accuracy
7. QQP GLUE paragraph detection Quora accuracy
8. RTE GLUE natural language inference news, Wikipedia accuracy
9. SST-2 GLUE sentiment analysis movie reviews accuracy
10. WiC SuperGLUE word sense disambiguation lexical databases accuracy
11. CB SuperGLUE natural language inference various accuracy
12. COPA SuperGLUE question and answering blogs, encyclopedia accuracy
13. BoolQA SuperGLUE boolean question and answering Wikipedia accuracy
14. MultiRC SuperGLUE question and answering various accuracy
15. IMDB SuperGLUE sentiment analysis movie reviews accuracy
TABLE XI: Four different orders of task sequences used for our experiments. Orders 1-2 correspond to the SuperNI benchmark. Orders 3-4 are long-sequence orders following [18].
Order Model Task Sequence
1 T5, LLaMA-2
task1572 → task363 → task1290 → task181 → task002 →
task1510 → task639 → task1729 → task073 → task1590 →
task748 → task511 → task591 → task1687 → task875
2 T5, LLaMA-2
task748 → task073 → task1590 → task639 → task1572 →
task1687 → task591 → task363 → task1510 → task1729 →
task181 → task511 → task002 → task1290 → task875
3 T5, LLaMA-2
mnli → cb → wic → copa → qqp → boolqa → rte → imdb →
yelp → amazon → sst-2 → dbpedia → ag → multirc → yahoo
4 T5, LLaMA-2
yelp → amazon → mnli → cb → copa → qqp → rte → imdb →
sst-2 → dbpedia → ag → yahoo → multirc → boolqa → wic

### v2

@@ VI-A1 Datasets
Additionally, we utilize the Long Sequence Benchmark [17], which consists of 15 classification datasets tailored for CL. Consistent with previous work [18], we randomly select 1,000 samples per task for training and reserve 500 samples per class for validation and testing. The experiments include two different task orders for each benchmark.
Further details about the tasks and their order are provided in the Appendix.
TABLE I: The overall results on two continual learning benchmarks with T5-large model. All results are averaged over two different orders of each benchmark.
@@ VI-A3 Baselines
We evaluate TaSL against the following PEFT-based CL baseline methods: SeqLoRA: sequentially trains the LoRA on the task orders. IncLoRA: incremental learning of new LoRA parameters on a sequential series of tasks. Replay: replays real samples from old tasks when learning new tasks to avoid forgetting. EWC [23]: finetune the model with a regularization loss designed to minimize updates to parameters critical to previously learned tasks. L2P [42]: uses the input to dynamically select and update prompts from a fixed prompt pool. LFPT5 [43]: continuously trains a soft prompt for each task with generative replay. Prog-Prompt [17]: sequentially concatenates previous learned prompts to the current one during the training and testing time. O-LoRA [18]: learns tasks in distinct LoRA subspaces, ensuring orthogonality, and aggregates all LoRA weights for testing. SAPT [15]: leverages pseudo samples and a shared attention framework to align PEFT block learning and selection.
Table II compares our TaSL with these baselines, highlighting three key advantages: data privacy-friendliness, model parameter efficiency, and improved generalization capabilities. To clarify the specific localization and consolidation techniques used in TaSL and its variants: TaSL: our foundational framework incorporates matrix-level skill units, utilizes a first-order gradient-based metric for skill localization (Eq. 8), and employs static weighted consolidation (Eq. 15). TasLoRA: as the most recent extension of TaSL, this variant introduces LoRA-tailored skill units, adopts a second-order gradient approximation metric for skill localization (Eq. 9), and features adaptive weighted consolidation (Eq. 17). TasL-M and TasLoRA-M: these enhancements of TaSL and TasLoRA integrate memory replay to further boost performance.
@@ VI-A4 Implementation Details
We utilize two distinct language model architectures previously adopted in the field of CL for NLP: the encoder-decoder T5 model [49] and the decoder-only LLaMA model [50]. In TaSL, the hyperparameters $\alpha_{1}$ and $\alpha_{2}$ in Eq. (10) and Eq. (11) are set to 0.85. We set $\beta$ in Eq. (13) to 0.7 and $\tau$ in Eq. (16) to 0.15. Following [51], we allocate 2% of the original training set volume for replay samples in all replay-based baseline methods. For different backbones, we utilized the following hyperparameters:
  • T5-base (220M), T5-large (780M), and T5-XL (3B): Training was conducted with a learning rate of 3e-4, batch size of 8, maximum input length of 512, maximum target length of 128, a rank of 8, targeting modules [q, v] and 10 epochs.
  • LLaMA (7B): With a learning rate of 3e-4, a batch size of 128, a cutoff length of 512, and 10 epochs. LoRA settings were rank = 8, alpha = 16, dropout = 0.05, targeting modules [q_proj, v_proj]. For testing, settings included temperature = 0.02, top_p = 0, top_k = 1, max new tokens = 128.
Experiments are conducted using 4 NVIDIA A100 GPUs.
@@ VI-B Main Results
The overall CL results of various methods using the same T5-large backbone are summarized in Table I.
Firstly, TaSL and its variants consistently demonstrate superior CL performance by effectively facilitating knowledge transfer. Compared to previous replay-free methods like IncLoRA and O-LoRA, our methods excel in addressing the critical challenges of CF and KT, achieving the highest AP, FWT, and BWT when learning tasks sequentially.
Furthermore, our enhanced version, TasLoRA, significantly outperforms TaSL. Specifically, TasLoRA achieves average gains of 1.3% in AP and 1.4% in BWT, underscoring the effectiveness of enhancements to each framework component. Notably, even without memory replay, TasLoRA nearly matches the performance of SAPT with memory replay on the Long Sequence Benchmark, particularly in BWT, with only a minor difference of 0.8% (-2.04% vs. -1.25%).
For memory-based methods, both TaSL-M and TasLoRA-M surpass the best SAPT models, further proving the versatility and robustness of our framework. For subsequent experimental analyses and comparisons, we will focus on TasLoRA, the most effective memory-free version, for further evaluations.
Secondly, TasLoRA consistently demonstrates superior performance across various backbones. To further validate our framework’s effectiveness, we conducted experiments using a range of parameter-level backbones. As depicted in Figure 4, performance improvements are evident with increasing model size. For instance, in T5-XL, TasLoRA boosts the AP metric from 28.6% to 33.3%, and shows substantial enhancements in FWT and BWT, rising from 11.1% to 16.1% and improving from -20.6% to -14.0%, respectively. These results further validate the robust generalization ability of our method.
@@
Figure 4: Performance of TasLoRA across various backbones.
Thirdly, our skill consolidation technique effectively counters catastrophic forgetting. To rigorously assess our model’s ability to mitigate forgetting, we evaluated its performance on the initial task after training on subsequent tasks using the Long Sequence Benchmark. Figure 5 illustrates that TasLoRA significantly reduces the rate of forgetting, with an average performance decline of only 10% after training on the final task. In contrast, vanilla backbones suffer a substantial average performance drop of 22%, highlighting our method’s enhanced capacity to preserve learned knowledge.
Figure 5: Performance trajectory of Task 1 on Long Sequence Benchmark throughout the continual learning process.
@@ VI-C Performance on Unseen Tasks
Building on prior research [15], we selected three tasks from each category to evaluate our method’s cross-task generalization capabilities, a critical measure for CL algorithms. The results, indicated by the average Rouge-L scores in Table III, highlight the effectiveness of our approach. T5-ZS refers to the zero-shot task adaptation methods employed. TasLoRA exhibited the best performance among all methods tested, a success largely due to its proficient identification and management of task-specific and shared parameters.
TABLE III: Results on unseen tasks using the T5-Large backbone.
@@ VI-D Visualization of Skill Units
We visualized the distribution of importance scores for skill units across tasks on the T5-large model, as shown in Figure 6 and Figure 7, leading to several critical insights:
  • There is noticeable variability in the importance of skill units for the same task, with important skill units comprising only a small fraction of all trainable LoRA parameters.
  • The distribution of important skill units is task-dependent, illustrating the presence of both task-shared and task-specific parameters. This observation substantiates the feasibility of the design motivations behind the TaSL framework.
  • For classification tasks, such as those in the Long Sequence Benchmark (Figure 6), the skill units in the encoder, particularly the lower layers closer to the model input, play a more pivotal role. Conversely, for generative tasks, like dialogue generation and summarization in the SuperNI benchmark (Figure 7), both encoder and decoder units are crucial, impacting both lower and upper layers of the network.
  • Within each layer, the importance of the Query (Q) matrices in the attention mechanism consistently exceeds that of the Value (V) matrices.
Figure 6: Visualization of importance scores for LoRA-tailored skill units across different tasks on T5-large within the Long Sequence Benchmark.
Figure 7: Visualization of importance scores for LoRA-tailored skill units across different tasks on T5-large within the SuperNI Benchmark.
TABLE IV: Ablation Study for Importance-Aware Skill Localization.
@@
VI-E Ablation Study
In this section, we assess the impact of importance-aware skill localization, fine-grained skill consolidation, and memory size on model performance. Detailed analysis of hyperparameter sensitivity is provided in the Appendix.
VI-E1 Effect of the Proposed Importance Metric in Skill Localization
We explore alternative importance scoring methods to validate the effectiveness of our approach: (i) To evaluate the impact of using moving averages on trajectory gradients, we modify $s(\cdot)$ in Eq. (6) to focus solely on sensitivity, as outlined in Eq. (9); (ii) To verify the validity of our proposed approximate second-order gradient importance metric in Eq. (6), we implement a first-order Taylor expansion as described in Eq. (8). The comparative results, presented in Table IV, show that using moving averages for importance scoring substantially enhances model performance. Additionally, compared to the first-order gradient approximation, our novel second-order gradient method results in performance improvements of up to 3.4%, 2.1%, and 4.2% across the evaluated metrics. These findings underscore the significant role of accurate skill localization in elevating model effectiveness.
@@ VI-E2 Effect of Adaptive Skill Consolidation
We evaluate our fine-grained, skill unit-level adaptive weighted consolidation against two coarse-grained strategies: (i) Weight-Ensemble, which uniformly averages the LoRA parameters using a global weight as per Eq. (14), and (ii) Exponential Moving Average (EMA) [52], which computes a running average of parameters at each fine-tuning iteration.
According to Table V, Weight-Ensemble significantly enhances the performance of the vanilla model, demonstrating the advantages of coarse-grained averaging in CL. While EMA generally outperforms Weight-Ensemble, it does not match the effectiveness of our fine-grained approach. EMA’s frequent parameter adjustments within the same task can lead to suboptimal outcomes. In contrast, our method, which strategically averages weights only after the completion of each task, boosts computational efficiency and model performance.
TABLE V: Ablation Study for Fine-Grained Skill Consolidation.
@@
VI-E3 The Effect of Memory Size
We examine how varying the memory size affects the performance of Replay, TaSL-M, and TasLoRA-M. By adjusting the memory size per task $|M|$ to 2%, 5%, 10%, 50%, we summarized the results in Table VI. As expected, increasing the memory size generally enhances the performance of all methods, with Replay showing the most significant improvement. This is predictable as Replay heavily depends on memory to mitigate catastrophic forgetting. When the memory size is substantially increased, Replay approaches the functionality of multi-task learning, which, while effective, incurs considerable storage and computational costs.
Conversely, TaSL-M and TasLoRA-M utilize skill consolidation strategies to effectively maintain parameters that store historical knowledge. Consequently, while increasing memory size does boost their performance, the improvements are less dramatic as the memory size grows.
TABLE VI: Ablation Study on Memory Size.
Memory Size
2% 5% 10% 50%
Replay 75.3 76.7 78.4 89.9
TaSL-M 84.2 85.1 85.7 91.2
TasLoRA-M 84.3 85.4 86.2 91.8
@@ VII Conclusion
In this paper, we introduced the novel Task Skill Localization and Consolidation (TaSL) framework for language model continual learning. TaSL employs importance-aware skill localization and fine-grained skill consolidation to effectively differentiate between task-specific and shared knowledge within each skill unit, mitigating forgetting while enhancing knowledge transfer. The framework is highly generalizable and extensible, allowing integration with memory replay methods to boost performance. Extensive testing shows that TaSL excels in preserving past knowledge and performing well in new tasks, surpassing previous state-of-the-art approaches. Comprehensive experiments demonstrate demonstrate our framework’s exceptional ability.
@@
Despite significant improvements in efficiency and performance in large language model continual learning, certain limitations persist. For instance, the choice of importance thresholds following the determination of importance distribution can impact the effectiveness of skill consolidation. Dynamically selecting these thresholds based on data distribution may yield a more precise classification of task-shared and task-specific parameters, further enhancing performance. Additionally, merging the skill localization and consolidation phases could allow for ongoing parameter consolidation based on their importance during model training. This integrated approach would facilitate more flexible adaptations and better address scenarios in online continual learning.